A man walks a certain distance at the rate of 5 miles an hour and returns at the rate of 4 miles an hour. If the total time that it takes him is 3 hours and 9 minutes, what is he total distance that he walks?
Answer by richwmiller(9144)

x+y=3 9/60;
5x=4y
x=7/5
y=7/4
7 miles